OpenStudy (asilva5762):

What is the value of (f-g)(144)

OpenStudy (quickstudent):

Welcome to Open Study, @asilva5762 ! :D

OpenStudy (asilva5762):

Hi 😄

OpenStudy (quickstudent):

So is this all there is to the expression, or does it equal something? in other words, are you solving for f and g, or are you simplifying the expression?

OpenStudy (asilva5762):

Expression

OpenStudy (quickstudent):

@asilva5762 , if you are simplifying the expression, you can do so by using the distributive property, which is this: (a)(b + c) = (a * b) + (a * c) So, (f - g)(144) = 144f - 144g

OpenStudy (asilva5762):

Thnks

OpenStudy (quickstudent):

Your Welcome! :)